OpenStudy (anonymous):

Write the product (5 sin 3pi/4)(cos 7pi/4) as a sum or difference. How would I do this problem? I know the answer, but I don't know how to do it. The answer is: 5/2 (sin 5pi/2 - sin pi).

OpenStudy (anonymous):

use 2sinacosb =sin (a+b) - sin(a-b)

OpenStudy (anonymous):

5/2 *2 sin 3pi/4)(cos 7pi/4

OpenStudy (anonymous):

here a= 7pi/4 and b = 3pi/4 thus a+b=5pi/2 and a-b = 4pi/4=pi

OpenStudy (anonymous):

Oh, okay. Thank you! :)

OpenStudy (raden):

it should 2sinacosb =sin (a+b) + sin(a-b) @matricked

OpenStudy (anonymous):

@RadEn yup u r correct dear but the above also works as b>a

OpenStudy (anonymous):

Thanks!!!! :D

OpenStudy (raden):

sorry, @matricked i didny see, ur coment that a= 7pi/4 and b = 3pi/4 :)

OpenStudy (anonymous):

no probs dear